### v3

# Split Unlearning

Abstract
We for the first time propose, implement, and evaluate a practical Split Unlearning framework by enabling SISA-based machine unlearning (SP’21) in Split Learning (SL).
We introduce SplitWiper and SplitWiper+, which leverage the inherent “Sharded” structure of SL and address the issues where existing SL methods compromise the “Isolated” principle of SISA due to the tight coupling between clients and the server. This coupling involves frequent bidirectional data flows and iterative training across all clients. We resolve these challenges by designing a new one-way-one-off propagation scheme. Our design decouples the propagation of neural signals between clients and between clients and the server, enabling SISA-based unlearning in SplitWiper, even with absent clients. SplitWiper+ further enhances client label privacy against the server under this scheme by using differential privacy. Our extensive experiments across diverse data distributions and tasks demonstrate that SplitWiper, which involves only the requesting clients, achieves 0% unlearning accuracy and is over 108% more effective in retaining accuracy than existing SL methods, while maintaining constant overhead through its one-way-one-off propagation, reducing computational and communication costs by over 99%. SplitWiper+ preserves over 90% of label privacy when sharing masked labels with servers.
1 Introduction
Split Learning (SL) is a key approach in privacy-preserving distributed learning, particularly in scenarios where clients with limited computational resources, such as smartphones or IoT devices, need to offload part of the training computations to servers [1, 2, 3]. Existing SL frameworks primarily include two configurations: Vanilla SL and U-shaped SL [4]. In Vanilla SL, the model is split into two segments: clients manage the segment receiving the training data and send intermediate values to the server, while the server handles the segment producing the final outputs without accessing raw data. U-shaped SL extends this by eliminating the need for label sharing, protecting both label privacy and data privacy. The model in U-shaped SL is divided into three segments: clients process the input data and produce output classes, while the server manages the intermediate layers.
Ensuring compliance with regulatory requirements for “data to be forgotten” (e.g., GDPR [5]), which grants individuals the right to delete their personal data, necessitates the development of machine unlearning [6] techniques to effectively remove specific data from models. However, the requirement becomes challenging within the context of SL, where the model segments across clients and servers need to be jointly updated. Current methods include retraining the model [7, 8] or using alternatives like fine-tuning, gradient ascent, and knowledge distillation [9, 10, 11]. While these alternatives are favored for their lower unlearning efficiency (i.e. time consumption) compared to retraining, which prioritizes effectiveness (i.e., accuracy[12], they all require the inclusion of all data samples during each update iteration. This becomes problematic when some data samples are absent due to contextual constraints. The Sharded, Isolated, Sliced, and Aggregated (SISA) method [6, 13, 14] addresses it by segmenting data and training multiple models in parallel, maintaining higher effectiveness and efficiency than non-SISA methods.
Research gaps. Existing SL frameworks [1, 2, 3] are inadequate for directly applying SISA in multi-client scenarios, where multiple clients have their independent models while sharing a common model on the server. Although the clients’ training data can be considered “Sharded”, SL undermines the “Isolated” requirement of SISA for two reasons: (1) the server model is tightly coupled with the client models through frequent bidirectional propagation, and (2) the shared server model is iteratively trained across all clients. Therefore, a new framework is necessary to decouple the shared server model from client models, enabling SISA-based unlearning in multi-client SL environments.
Contributions. In this paper, we address these gaps in enabling SISA-based unlearning in multi-client SL environments where unlearning can thus be conducted smoothly even with absent clients. To be specific:
[…]
2 Preliminaries
2.1 Split Learning
Key factors. SL is a distributed machine learning technique designed to train neural networks without requiring the sharing of raw data between $[1,K]$ clients, each owning datasets $D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$, and servers that hold no data [4, 15]. A typical SL solution involves two key factors: topology and training.
  • $\bullet$
    Topology. The network is divided into two parts, referred to as client-side and server-side. We denote the network as a sequence of layers $\{L_{1},L_{2},\ldots,L_{n}\}$. The topology step involves splitting this sequence into two subsets: the client-side layers for $K$ clients in parallel are denoted as $\{\{L_{1}^{1},L_{2}^{1},\ldots,L_{m}^{1}\}$, $\ldots$, $\{L_{1}^{K},L_{2}^{K},\ldots,L_{m}^{K}\}\}$ and the server-side layers $\{L_{m+1},L_{m+2},\ldots,L_{n}\}$. The clients and server initialize their respective portions of the network randomly.
  • $\bullet$
    Training. Each client-$k$ processes its local data through the initial layers of the model, producing an intermediate output $H_{k}^{(t)}=\mathcal{F}_{o}^{k}(X_{k};W_{k}^{(t)})$ through its layers $\{L_{1}^{k},L_{2}^{k},\ldots,L_{m}^{k}\}$, where $W_{k}^{(t)}$ represents client-$k$’s weights at iteration $t$. This output is sent to the server, which continues the forward propagation through its layers $\{L_{m+1},L_{m+2},\ldots,L_{n}\}$ and updates its own weights $W_{s}^{(t)}$ based on the gradients: $W_{s}^{(t+1)}=W_{s}^{(t)}-\eta\Phi\left(\frac{\partial\mathcal{L}_{1}}{\partial W_{s}^{(t)}},\frac{\partial\mathcal{L}_{2}}{\partial W_{s}^{(t)}},\dots,\frac{\partial\mathcal{L}_{K}}{\partial W_{s}^{(t)}}\right)$, where $\frac{\partial\mathcal{L}_{k}}{\partial W_{s}^{(t)}}$ is the gradient contribution of loss values from client-$k$ and $\Phi(\cdot)$ is a customized aggregation function. The server then sends back the gradients to each client, allowing them to update their own weights: $W_{k}^{(t+1)}=W_{k}^{(t)}-\eta\frac{\partial\mathcal{L}}{\partial W_{k}^{(t)}}$.
The processes described above enable training without the need to share raw data or specific model details between the client and server. In scenarios involving multiple clients and a single server, two primary training methods are commonly employed. The first method synchronizes model weights between clients after each epoch, while the second method operates without synchronization, with clients alternating epochs in collaboration with the server [4].
Incremental training process. This process has been utilized by many prevalent learning frameworks including Federated Learning (FL) [16, 17] and SL, where a model is continuously updated by gradually incorporating new data over time. Each iteration of back-propagation needs to be sent back to every client involved in the training process and builds incrementally on the previous ones, embodying the incremental nature of SL [18]. This process enables the model to dynamically learn and adapt as new data becomes available, ensuring it remains responsive to evolving patterns and trends.
The challenges of applying SISA in multi-client SL scenarios stem from the incremental training process inherent in these setups. This iterative process across multiple clients undermines the Isolated requirement of SISA due to the close integration of the server and client model segments through regular forward and backward propagations, and the shared server model segment being trained incrementally across all clients. The knowledge to be unlearned, reflected in $W_{k}^{(t)}$, would have also influenced $W_{k^{\prime}}^{(t^{\prime})}$ for all $k^{\prime}\neq k$ and $t^{\prime}\geq t$, necessitating the recalculation of corresponding changes in subsequent updates, thereby making it impossible to apply SISA unlearning [19, 12].
2.2 Machine Unlearning
Machine unlearning involves removing the influence of specific data subsets from a trained model while preserving its performance on the remaining data [10, 9]. The goal is to eliminate the impact of certain labels on the model’s predictions without compromising accuracy on the retained data. Let $D=\{\left(X_{i},Y_{q}\right)\}_{i=1,q=1}$ represent the dataset, where $X_{i}$ is the $i$-th training sample and $Y_{q}\in\{1,\cdots,Q\}$ is its label. Unlearning removes the influence of subset $D_{u}\subset D$ with labels $Y_{u}$ while preserving performance on the remaining data $D_{r}=D\backslash D_{u}$ with labels $Y_{o}$, ensuring $Y_{u}$ is forgotten and predictions on $Y_{o}$ remain accurate.
Retraining. The most direct and effective, though not efficient, method of implementing unlearning is retraining [7, 8]. In this method, the revoked sample is deleted, and the model is retrained on the original dataset minus the deleted sample, denoted as $D_{r}=D\setminus{((X_{i},Y_{q})}\subset D_{u})$. While this method is effective and straightforward, the computational overhead becomes prohibitive for complex models and large datasets.
Alternatives. Referring to fine-tuning, gradient ascent, and knowledge distillation, These unlearning methods are proposed to improve unlearning efficiency compared to retraining, but they still consider retraining as the baseline for unlearning effectiveness [9, 10, 11], including:
  • $\bullet$
    Fine-tuning. Adjusting the model parameters by further training on $D_{r}$ to deliberately induce catastrophic forgetting of $D_{u}$, typically using $W^{(t+1)}=W^{(t)}-\eta\nabla\mathcal{L}_{r}(W^{(t)})$, where $\mathcal{L}_{r}$ is the loss function computed on the retained data.
  • $\bullet$
    Gradient ascent. Updating the weights in the direction of the gradient of a loss function to maximize the objective: $W^{(t+1)}=W^{(t)}+\eta\|\nabla\mathcal{L}(W^{(t)})-\nabla\mathcal{L}_{r}(W^{(t)})\|$.
  • $\bullet$
    Knowledge distillation. Trading-off between a preserver that aids to retain knowledge and a destroyer that aids in forgetting knowledge: $\mathcal{L}=\alpha\mathcal{L}_{u}+(1-\alpha)\mathcal{L}_{r}$.
SISA. SISA [6] is a guideline designed to enable efficient and independent unlearning for those who request it, without affecting other participants. In SISA, the original training set $D_{o}$ is divided into $K$ disjoint shards {$D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{K}$}, each used to train a separate model $\mathcal{F}_{o}^{1},\mathcal{F}_{o}^{2},\ldots,\mathcal{F}_{o}^{K}$}. These models generate individual predictions, which are then aggregated to produce a global prediction. When an unlearning request is made, only the shard containing the specific data sample needs to be retrained, allowing the unlearning process to be conducted independently for the requesting party, without impacting other clients or the overall system.
Why SISA-based SL. Given the unique challenges of multi-client SL, the ability to perform machine unlearning with absent clients due to network issues is essential. While retraining and other alternatives may focus on either effectiveness or efficiency, SISA enhances these methods by enabling independent unlearning for requesting clients without affecting others. In SISA, any unlearning method can be applied within each division, converting non-SISA structures into SISA-based ones. This capability is crucial in SL, making SISA the most viable method for addressing the limitations of current non-SISA methods in SL. For simplicity, this paper assumes retraining is applied on each client.
3 SISA-based Split Unlearning
3.1 Problem Definition
In the context of SL, the training set $D_{o}$ is distributed among $K$ clients, and a server that serves as a training service provider. In the rest of this paper, we consider that the dataset is distributed in a horizontally non-IID setting of data distribution, i.e., each client owns a subset of $D_{o}$ with different feature distributions and uneven label distributions.
Multi-client settings. We focus on multi-client scenarios. For a trained model $\mathcal{F}_{o}$, we can divide it into a client-side part $\mathcal{F}_{o}^{k}$ for client-$k$, and a server-side part $\mathcal{F}_{o}^{s}$. Each client-$k$ owns its $\mathcal{F}_{o}^{k}$ in parallel and shares a single $\mathcal{F}_{o}^{s}$. We define the evaluation of model performance from the perspective of client-$k$ as $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, where $\mathcal{T}$ represents the evaluation function. Each client functions as an individual user managing its local data, aiming to enhance the prediction accuracy of its pre-trained local model by leveraging the server’s computational capabilities. Consequently, the performance assessment focuses on the server’s perspective, which aggregates the independent evaluations, $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k}),\forall k$.
Independent unlearning. Accordingly, any client-$k$ that raises an unlearning request should be conducted only between this particular client and the server, with no needs of other $(k^{\prime}\neq k)$-th clients being involved in the unlearning process. This means the server should unlearn client-$k$’s unlearned features from the server-side part $\mathcal{F}_{o}^{s}$ upon the intermediate outputs of the updated client-side part $\mathcal{F}_{o}^{k}$. Formally, to respond to the unlearning request for a series of data samples $D_{u}^{k}$ raised by client-$k$, the entire SL network needs to satisfy an unlearned model $\mathcal{F}_{u}$ trained on $D_{u}=D_{o}\setminus\{D_{u}^{k}\}$.
Challenges of SISA-based split unlearning. The split unlearning process faces what we term the Dilemma of Independence and Effectiveness, particularly when clients operate independently with limited training resources. Non-SISA methods, involving all clients, are vulnerable to client absence during unlearning and impose significant overhead on both the initiating client ($k$) and other clients ($k^{\prime}\neq k$). This method unfairly forces disinterested clients-$k^{\prime}$ to expend resources solely to meet client-$k$’s needs.
Figure 1: Architectural Design. In SplitWiper (Sec.3), each client-$k$ trains its local model $\mathcal{F}_{o}^{k}$ on its dataset $D_{o}^{k}$, treating them as SISA shards without additional partitioning. After training, clients freeze the weights and send outputs with labels to the server for caching (one-off), while the server continues training $\mathcal{F}_{o}^{s}$ using cached values without returning to the clients (one-way). SplitWiper+ (Sec.4) enhances privacy by using a label expansion strategy to convert real labels into masked ones, which clients then share with the server for label-protected training.
The pursuit of SISA-based split unlearning is to devise an unlearning algorithm fulfilling the following goals:
  • $\bullet$
    G1: Independent unlearning. This goal emphasizes the importance of performing the unlearning operation without affecting other clients ($k^{\prime}\neq k$) that are not involved in the unlearning process. This ensures that the unlearning is confined to the requesting client and server, leaving the operations and resources of other clients untouched.
  • $\bullet$
    G2: Effective unlearning with retained utility. This goal focuses on ensuring that the unlearning process thoroughly removes the influence of the unlearned data from the model, while also maintaining or enhancing the predictive accuracy for the retained data.
3.2 SplitWiper
We propose SplitWiper, which enables SISA-based unlearning. The Sharded training of SplitWiper involves multiple clients, each acting as a distinct shard with different data samples. Our one-way-one-off propagation scheme ensures Isolated training by eliminating backward gradient propagation (one-way) and sending each client’s intermediate output to the server only once (one-off). This confines the impact of each shard to its corresponding client, with no information shared between different clients’ models. Within each client, data can be further divided into Slices by embedding the SISA scheme. On the server side, Aggregation combines cached intermediate outputs from clients, refining the server’s model and enabling efficient unlearning across multi-client SL.
3.2.1 Learning in SplitWiper
The SplitWiper framework (cf. Fig. 1) consists of the following three learning phases (also see Algorithm 1): client model pre-training, client model freezing and caching, and server model training.
Client model pre-training (Line 6 - Line 8). The concept of SISA is applied in this phase in a sense that the non-IID distributed datasets {$D_{o}^{1},D_{o}^{2},\ldots,D_{o}^{k}$} owned by any client-$k$ are inherently treated as the shards of SISA. By default, SplitWiper does not involve shard partitioning, as the server lacks access to any client’s dataset and only provides training services for outsourcing tasks. Consequently, each client-$k$ pre-trains its local model $\mathcal{F}_{o}^{k}$ on its own dataset $D_{o}^{k}$ independently and in parallel. This setup allows clients to use entirely different datasets and training tasks, as long as the dimensions of their outputs match the server’s input requirements to ensure smooth integration and communication.
Client model freezing and caching (Line 9). The new one-way-one-off propagation is crucial by freezing clients’ weights and caching them on the server. After client model training, the outputs of the last layer (the cut or intermediate layer) and corresponding labels are shared with the server for caching, while weights are frozen to prevent further computation and communication during server training. In SplitWiper, mechanisms such as dropout layers or batch normalization are managed solely by the server, ensuring client model determinism. This allows for consistent sharing of intermediate outputs, which the server can cache without concerns about client model variability.
Server model training (Line 15 - Line 17). As the central server that hosts the final output layer of the entire network, the server’s training process for its model, $\mathcal{F}_{o}^{s}$, functions similarly to the aggregation function $\Phi(\cdot)$ in SISA. By receiving and storing intermediate outputs from each client, the server accelerates the training of its model, which is typically much larger than the client models. With the client-side weights frozen, gradients are no longer back-propagated to the clients, restricting the incremental training process to the server. This ensures that subsequent updates do not propagate to all clients, implementing the one-way-one-off propagation scheme. This method lays a strong foundation for enabling unlearning with absent clients while significantly reducing overhead.
3.2.2 Unlearning in SplitWiper
Our G1 ensures that only the clients requesting unlearning need to participate, while others remain uninvolved. We give below the workflow of SISA-based split unlearning with SplitWiper (cf. Algorithm 2 in Appendix) The algorithm takes as input the number of epochs for client model updating $N$, the number of epochs for server model updating $M$, and the initiating client-$k$ who proposes the requirement for unlearning, and outputs the updated predicted label on the $Concat(\mathcal{F}_{u}^{k},\mathcal{F}_{u}^{s})$. It works in three steps as follows:
  • $\bullet$
    Step 1: Client weight unfreezing and dataset modification (Line 4 - Line 6). In this step, client-$k$ that initiates the retraining request unfreezes its weights and removes the intended unlearned samples from its dataset, $D_{r}^{k}=D_{o}^{k}\setminus\{D_{u}^{k}\}$. This prepares the local model for unlearning.
  • $\bullet$
    Step 2: Client model unlearning and freezing (Line 7 - Line 10). The initiating client-$k$ retrains its model for $N$ epochs and then freezes the updated weights. Subsequently, it sends intermediate outputs and corresponding labels to the server for further training in the server.
  • $\bullet$
    Step 3: Server model updating (Line 13 - Line 17). The server updates its cache with client-$k$’s intermediate outputs, trains the server model for $M$ epochs using the cached values, and then obtains the updated output distribution.
Given that G2 (effective unlearning with retained utility) is a critical requirement for any feasible unlearning method, this unlearning process achieves G1 (independent unlearning) by having clients freeze their weights and cache intermediate values on the server side, selectively unfreezing them at the appropriate time to facilitate the unlearning process without involving or disrupting others.
4 SplitWiper+: An Privacy-Preserving Variation via Label Expansion
The one-way-one-off propagation halts learning propagation at the server side and freezes the client’s parameters after the clients’ outputs are cached for the first time. This prevents the need to trace back to clients for label matching, which is, however, a necessity when labels cannot be shared with the server in U-shaped SL. This introduces a key challenge if keeping labels private is indeed a requirement:
How to keep labels private while still enabling their sharing with the server for the unlearning process, without compromising the integrity of G1 and G2?
Player model. Two types of players are considered in this label-privacy-preserving scenario.
  • $\bullet$
    Internal users. This includes clients and servers involved in the training process. They are honest-but-curious [20], where they follow the prescribed protocol without tampering with the training process by sending incorrect features or causing intentional failures. Each client is aware of the shared labels among all participants but prefers to keep their complete set of labels private from the others. Clients aim to conceal the quantity and the semantic information of labels from the server. This player model focuses on the level of privacy preservation in internal label sharing.
  • $\bullet$
    External attackers. An external attacker is considered to know the distribution of the original dataset fed into the input layer at the targeted client, and the posterior distribution of the model at the server side. The attacker aims to use a shadow dataset that mimics these distributions to construct an attack model for conducting inference attacks [21]. This player model focuses on the level of privacy preservation in the unlearning capability of the framework.
Considering the player model and keeping labels private while achieving efficiency and effectiveness, we introduce an additional objective beyond G1 and G2 as follows.
  • $\bullet$
    G3: Privacy-preserving label sharing. This goal emphasizes that the learning and unlearning processes should not disclose any label information from the client side to the server, including the quantity and semantics of the labels.
To achieve this, we introduce a label expansion strategy combined with a DP-based masking scheme that protects the semantic information and quantity of labels shared with the server during SL. This method is crucial when the server’s trustworthiness is uncertain, as it conceals real labels, protecting client privacy and data security. By effectively masking both the quantity and specifics of the labels, clients can securely transmit masked labels to the server, mitigating privacy and security concerns.
Figure 2: Label Expansion. Our proposed strategy preserves label quantity and semantics by expanding, shuffling, and anonymizing them, with intermediate values expanded via a DP mechanism.
Compared to U-shaped SL tailored for label privacy, our new DP-based method in SplitWiper+ requires only one-way communication between clients and the server. This significantly reduces the communication frequency between clients and the server, enhancing efficiency and lowering operational costs. The process of the proposed label expansion-and-masking scheme is illustrated in Fig. 2.
4.1 Label Expansion Strategy
We consider a scenario with $Q$ real labels, represented as $\mathbb{Y}={Y_{1},\cdots,Y_{Q}}$, where each client-$k$ holds a subset $\mathbb{Y}_{k}={Y_{1},\cdots,Y_{Q_{k}}}$. These subsets collectively form the entire label set, such that $\bigcup_{k=1}^{K}\mathbb{Y}_{k}=\mathbb{Y}$. The corresponding intermediate values for label $Y_{q}$ on client-$k$ are denoted as $V_{q}^{k}$. In Vanilla SL, where label privacy is not a concern, the data exchanged between client-$k$ and the server consists of pairs of labels and their corresponding intermediate values $\{Y_{q},V_{q}^{k}\}$. To address the risks of label leakage, SplitWiper+ modifies this method by sharing masked labels and their DP-protected intermediate values $\{Y_{q}^{*},{V_{q}^{*}}^{k}\}$ instead.
To address the potential risk of sensitive label information being leaked during label sharing, we introduce a method that allows for label sharing without revealing the full set of labels $\mathbb{Y}_{k}$ held by client-$k$ (cf. Algorithm 3 in Appendix). Given the non-IID data distribution across clients, implementing a unified label expansion strategy in SplitWiper+ requires additional communication among clients. This strategy only requires clients to agree on all labels used in the SL task, without the need to exchange the actual data samples. Since internal users are assumed to be honest but curious, direct label sharing among clients could still expose sensitive information, making our method crucial for preserving privacy.
Secure label consensus building (Line 5 - Line 12). The communication among clients facilitates consensus among all clients on the complete set of real labels $\mathbb{Y}$ by allowing them to sequentially contribute their labels to a communal pool. The process is governed by two probabilistic thresholds:
  • $\bullet$
    $\lambda_{1}$: The probability that client-$k$ will select and add labels from its set $\mathbb{Y}_{k}$ that are absent in the communal pool $\mathbb{Y}$.
  • $\bullet$
    $\lambda_{2}$: The probability that client-$k$ will continue to participate honestly in the sharing process after assessing whether $\mathbb{Y}_{k}$ is already included in $\mathbb{Y}$.
These probabilistic controls over label sharing and continued participation ensure that adversaries cannot deduce the complete label set $\mathbb{Y}_{k}$ of any individual client based on publicly shared information during the process. Even if an untrustworthy client leaks information about this process, the adversary, potentially the curious server, remains unaware of the specific labels held by each client.
Label expansion and anonymization (Line 13 - Line 16). Once the complete set $\mathbb{Y}$ is aggregated, clients collaborate to assign an expansion factor $\gamma_{q}$ to each label $Y_{q}$. This factor, $\gamma_{q}$, can be a randomly selected integer within a predefined range $\Gamma_{p}$ for each $Y_{q}$, or be a constant integer $\gamma>1$ and uniformly applied across all labels, i.e., $\gamma_{q}=\gamma,\forall q$. Each label, $Y_{q}$, is then expanded to $\gamma_{q}$ instances, forming the initial expanded label set $\mathbb{Y}_{exp}=\{Y_{1}^{*},\cdots,Y_{e}^{*},\cdots Y_{E}^{*}\}$, where $E=\sum_{q=1}^{Q}{\gamma_{q}}$. This expansion process masks the total count $Q$ of real labels, preventing the server from inferring specific client tasks or data types based on label quantity. The elements within $\mathbb{Y}_{exp}$ are shuffled and re-indexed to anonymize label semantics. While we use a basic pseudonymization technique, our framework can also accommodate methods like hashing [22] or encryption [23] based on specific needs. The mapping between $\mathbb{Y}$ and $\mathbb{Y}_{exp}$ is preserved in $\mathcal{G}_{Y}$, ensuring that labels are expanded and anonymized without losing semantic integrity.
Since the information shared between client-$k$ and the server consists of pairs of labels and their corresponding intermediate values $\{Y_{q},V_{q}^{k}\}$, when $Y_{q}$ is expanded into $\gamma_{q}$ masked labels, the corresponding $V_{q}^{k}$ also need to be expanded to match the masked labels. During this expansion, we introduce a DP mechanism to obfuscate $V_{q}^{k}$ for further privacy protection (cf. Algorithm 4 in Appendix).
DP-based expansion-and-masking scheme (Line 6 - Line 10). Following the pre-determined label expansion map $\mathcal{G}_{Y}$, Client-$k$ takes the pairs $\{Y_{q},V_{q}^{k}\}$ obtained from its local model $\mathcal{F}_{o}^{k}$ as input. Each $Y_{q}$ is then expanded into $\gamma_{q}$ masked labels, and $V_{q}^{k}$ is correspondingly expanded into $\gamma_{q}$ DP-protected intermediate values, resulting in the set $\mathbb{U}_{exp}=\bigcup\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$. Let the total collection of DP-protected intermediate values across all clients as $\mathbb{V}_{exp}=\{{V_{1}^{*}},\cdots,{V_{e}^{*}},\cdots,{V_{E}^{*}}\}$, keeping consistency with $\mathbb{Y}_{exp}$. The expansion of the intermediate values is carefully documented in $\mathcal{G}_{V}$, establishing a mapping that pairs expanded labels with their corresponding intermediate values, so we have $(\mathcal{G}_{Y},\mathcal{G}_{V})$ satisfying:
$(\mathcal{G}_{Y},\mathcal{G}_{V}):\{Y_{q},V_{q}\}\rightarrow\bigcup\{Y_{e}^{*},{V_{e}^{*}}\}.$ (1)
The mapping $(\mathcal{G}_{Y},\mathcal{G}_{V})$ represents a one-to-many expansion. Without collusion between clients and the server, $(\mathcal{G}_{Y},\mathcal{G}_{V})$ remains confidential, preventing the server from knowing the actual quantity or specific content of labels and intermediate values used in the training task.
4.2 Masked Label Sharing in SplitWiper+
Masked label sharing. In SplitWiper+, after freezing the pre-trained model, clients apply an expansion-and-masking scheme before sharing their data with the server. In contrast, SplitWiper involves clients directly sharing the obtained ${Y_{q},V_{q}^{k}}$ with the server after freezing the pre-trained model, without the additional masking step. By transforming $\{Y_{q},V_{q}^{k}\}$ into $\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$, clients transmit the processed masked labels and their corresponding DP-protected intermediate values to the server, without revealing the original information.
Theorem 1 (cf. Appendix A for proof) establishes that with the DP protection, the server cannot discern whether any two received labels are expanded from the same real label based on the corresponding intermediate values. In other words, the server cannot infer the quantity and content of real labels based on the received information, ensuring the privacy protection of the genuine labels used by the client.
Theorem 1.
Consider that $\mathbb{V}_{exp}=\{{V_{1}^{*}},{V_{2}^{*}},\cdots,{V_{E}^{*}}\}$ is expanded by $\mathbb{V}=\{{V}_{1},{V}_{2},\cdots,{V}_{Q}\}$ using a given method $\mathcal{G}_{V}$, where $\forall{V}_{q}\in\mathbb{V}$, $\gamma_{q}$ copies with $\varepsilon$-DP-based noise are generated as $\{{V_{a_{1}}^{*}},\cdots,{V_{a_{\gamma_{s}}}^{*}}\}\in\mathbb{V}_{exp}$. The probability that any two elements in $\mathbb{V}_{exp}$ correspond to the same element in $\mathbb{V}$ satisfies $2\varepsilon$-DP.
Label demasking. The server in SplitWiper+ proceeds with the next stage of training based on the received masked labels and intermediate values. The final training output on the server should yield masked predicted labels, which does not hinder the server-side training since it utilizes the same masked labels for learning.
We introduce an additional label demasking step on the client side when a client needs to utilize the complete model, i.e., $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$. This step converts the predicted masked labels from the server back into their original, unmasked forms to fulfill the task at hand. Given that $\mathcal{G}_{Y}$ is a one-to-many expansion method, its inverse, $\mathcal{G}_{Y}^{-1}$, is a many-to-one function as defined:
$Y_{q}=\mathcal{G}_{Y}^{-1}(Y_{e}^{*}).$ (2)
Each masked label corresponds directly to a real label, and since the mapping $\mathcal{G}_{Y}$ is known to all clients, they can effortlessly demask the predicted results received from the server by applying the inverse mapping $\mathcal{G}_{Y}^{-1}$. This demasking process is not merely a reversal but an integral part of maintaining the integrity and applicability of the model’s predictions. The demasked labels are then used to compute the performance $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, ensuring that the model evaluation is based on accurate and relevant data, reflective of true model performance in operational settings.
Label protection in the unlearning process. When the initiating client-$k$ requests unlearning, only client-$k$ participates in the subsequent update process. Other clients do not receive any additional updated information. The server, upon receiving the updated masked labels and corresponding DP-protected intermediate values shared by client-$k$, gains insight into the aspects of the unlearning process. However, since the labels are masked, the server cannot accurately infer the quantity or category of the real labels involved in the unlearning from the changes in the number of masked labels. Thus, we deem that the privacy associated with the unlearning labels remains protected, effectively safeguarding sensitive information from inadvertent disclosure.
Natural barrier against shadow inference. Shadow inference attacks involve an adversary using a shadow dataset to mimic the training and unlearning processes of the target model, ultimately allowing them to infer which data points have been unlearned, thereby compromising the privacy of the dataset [21]. The proposed label expansion strategy addresses this vulnerability and strengthens client privacy by implementing two key methods:
  • $\bullet$
    Training on private datasets. When training on private datasets, an external attacker, even with partial access to a victim’s data, cannot construct a comprehensive shadow dataset. Although [21] assumes the attacker knows all labels in the victim’s dataset, this is unrealistic unless using a well-known public dataset. The need to balance positive and negative cases and diversify shadow models is crucial for generalization in the attack model. Without full label knowledge, inferring the complete input-output mapping becomes difficult, limiting the attacker’s ability to replicate the training environment and conduct effective inference.
  • $\bullet$
    Training on various tasks. When clients are engaged in different tasks, such as CIFAR-10/MNIST, the attacker’s challenge increases, as they cannot easily reconstruct a shadow dataset that accurately reflects the diverse tasks of the clients. This further diminishes the attacker’s ability to predict the model’s behavior and the impact of unlearning.
For simplicity in our experiments (cf. Sec.5.1), we use public datasets to mimic this process and intentionally ignore the potential risk of inference posed by using public datasets.
SplitWiper as a specific case of SplitWiper+. The label expansion strategy provides inactive privacy protection when expansion factor $\gamma_{q}=\gamma=1,\forall{q}$, in which case SplitWiper+ becomes SplitWiper with no DP applied. For ease of comparison, throughout this paper, we will refer to SplitWiper as the case where $\gamma=1$ and SplitWiper+ as the case where $\gamma>1$ is applied. These two will be differentiated whenever label privacy is relevant in the experiments; otherwise, SplitWiper will be used by default.
Compared to SplitWiper, the label expansion strategy in SplitWiper+ introduces additional computational overhead and increases the size of the data transmitted from clients to the server, thereby increasing communication costs. Thanks to the one-way-one-off propagation scheme, these additional overheads are incurred only once, preserving the significant advantages over existing SL methods. Table 1 presents a comparison of the complexities among SplitWiper, SplitWiper+, and existing Vanilla and U-shaped SL. A detailed analysis of complexities can be found in Appendix B.
Table 1: Client-side Privacy and Overhead in SL
Framework Label Privacy Computational Communication
Learning Vanilla SL Low $O(M\cdot K\cdot\mathcal{H}_{K})$ $O(M\cdot K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$
U-shaped SL High $O(M\cdot K\cdot(\mathcal{H}_{K}+\mathcal{H}_{out}))$ $O(M\cdot K\cdot\mathcal{I}_{K})$
SplitWiper Low $O(K\cdot\mathcal{H}_{K})$ $O(K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$
SplitWiper+ Medium $O(K\cdot(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp}))$ $O(K\cdot\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$
Unlearning Vanilla SL Low $O(M\cdot K\cdot\mathcal{H}_{K})$ $O(M\cdot K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$
U-shaped SL High $O(M\cdot K\cdot(\mathcal{H}_{K}+\mathcal{H}_{out}))$ $O(M\cdot K\cdot\mathcal{I}_{K})$
SplitWiper Low $O(\mathcal{H}_{K})$ $O(\mathcal{I}_{K}+\mathcal{I}_{Y})$
SplitWiper+ Medium $O(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp})$ $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$
  • High: no label sharing; Low: real label sharing; Medium: masked label sharing; $\mathcal{H}_{K}$: client-side training complexity; $\mathcal{H}_{exp}$: label expansion-and-masking scheme complexity; $\mathcal{H}_{out}$: output layer complexity; $\mathcal{I}_{K}$: size of intermediate outputs; $\mathcal{I}_{Y}$: size of labels ; $\mathcal{H}_{out}$: output layer complexity.
5 Evaluation
In this section, we empirically evaluate SplitWiper of SplitWiper+, examining its compliance with key objectives: G1 (independent unlearning), G2 (effective unlearning with retained utility), and G3 (privacy-preserving label sharing).
5.1 Experimental Settings
[…]
SplitWiper and SplitWiper+ are versatile, accommodating a variety of unlearning scenarios including class-, client-, and sample-level tasks. The experiments specifically focus on class-level unlearning tasks for clarity of presentation [24], where $\mathbb{C}_{u}$ selects a subset of label, denoted as $\mathbb{Y}_{u}$ for unlearning, aiming to effectively erase all related information without affecting other clients $\mathbb{C}_{o}$, or other labels $\mathbb{Y}_{o}$. We set $\lvert\mathbb{Y}_{u}\rvert=1$, i.e., only a specific label is required for unlearning. The unlearned label set, $\mathbb{Y}_{u}$, is exclusively possessed by $\mathbb{C}_{u}$, aligned with the non-IID data distribution setting and enabling a clearer demonstration of unlearning effectiveness.
Dataset and data distribution. The experiments are conducted under various settings based on the dataset used and the degree of non-IID distribution among the clients:
  • $\bullet$
    Clients with same task. Clients operate on the same dataset but under different conditions:
    • $\circ$
      Homogeneous labels but heterogeneous volumes. All clients possess every label except $Y_{u}$, but the volume of labels varies among clients.
    • $\circ$
      Partially overlapping labels. Clients have different types and quantities of labels with some overlap, i.e., all labels except $Y_{u}$ are randomly distributed among 1-3 clients.
    • $\circ$
      Heterogeneous labels. Each client has a unique set of labels, with no overlap.
  • $\bullet$
    Clients with different tasks. Clients utilize different datasets for separate tasks, with completely distinct labels among the clients.
In each setting, the training and testing datasets are distributed among the clients using the same method to ensure consistency between the clients’ training and testing data.
For experiments with clients performing the same task, we use CIFAR-10 [25] and CIFAR-100 [26] datasets, which are widely utilized in the fields of computer vision and neural networks. CIFAR-10 is a collection of $60,000$ $32\times 32$ color images across $10$ classes, and CIFAR-100 includes 100 classes with $600$ images each. In CIFAR-10, we designate the label $Y_{u}$ as “bird”, and in CIFAR-100, as “aquarium-fish”. As for clients with different tasks, $3$ clients use CIFAR-10 and $2$ clients use MNIST [27] that includes a dataset of $60,000$ $28\times 28$ single-channel handwritten digit images. In this setup, “bird” from CIFAR-10 remained as $Y_{u}$. These diverse experimental setups help illuminate the strengths and challenges of implementing SplitWiper across different scenarios and datasets, reflecting its effectiveness in managing complex unlearning demands in non-IID distributed SL environments.
Table 2: Evaluation on Efficiency and Effectiveness over VGG across Different SL Frameworks
(Client-side) Training / Transmission Time (s) Accuracy (%)
Existing SL SplitWiper SplitWiper+ Existing SL SplitWiper SplitWiper+
U-shaped Vanilla $\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$ U-shaped Vanilla $\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$
Clients with same task Homo labels Hetero volumes CIFAR-10 L. $\mathbb{C}_{u}$ 74.36/216.40 66.03/261.63 2.23/9.04 9.73/18.08 15.85/27.13 9.08/18.08 $\mathbb{Y}_{u}$ 95.40 92.70 99.80 98.10 98.70 97.40 97.90 99.40 99.70
$\mathbb{C}_{o}$ 58.88/172.98 42.65/168.02 1.42/5.90 6.16/11.80 10.25/17.70 5.72/5.90 $\mathbb{Y}_{o}$ 93.47 92.36 97.04 96.67 97.41 97.07 96.16 96.83 97.97
U. $\mathbb{C}_{u}$ 34.23/168.63 22.18/168.72 1.43/5.84 5.94/11.68 10.20/17.52 5.64/5.84 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 57.48/172.95 42.89/168.42 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 94.23 92.94 97.22 96.47 97.91 96.78 97.07 97.14 98.56
CIFAR-100 L. $\mathbb{C}_{u}$ 79.88/324.58 74.68/328.02 1.44/8.24 6.31/16.48 10.70/24.73 5.88/24.72 $\mathbb{Y}_{u}$ 91.20 92.30 97.60 97.00 97.10 97.00 98.10 97.40 97.80
$\mathbb{C}_{o}$ 77.21/312.58 71.84/315.24 1.41/6.22 5.98/12.44 10.21/18.66 5.48/6.22 $\mathbb{Y}_{o}$ 82.01 81.77 85.39 83.75 84.17 82.53 83.44 85.27 86.08
U. $\mathbb{C}_{u}$ 77.03/308.95 70.93/312.04 1.36/6.13 5.92/12.25 10.17/18.38 5.43/6.13 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 77.15/309.45 71.56/312.04 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 82.79 82.02 86.86 83.51 84.95 82.71 84.23 86.19 87.17
Partially overlapping labels CIFAR-10 L. $\mathbb{C}_{u}$ 87.58/313.31 76.24/314.44 2.47/10.72 11.07/21.44 18.93/32.17 9.84/21.44 $\mathbb{Y}_{u}$ 95.60 94.30 99.60 99.50 98.40 99.90 99.20 99.60 99.70
$\mathbb{C}_{o}$ 42.99/162.61 38.59/157.22 1.28/5.34 5.53/10.68 9.43/16.02 5.47/10.42 $\mathbb{Y}_{o}$ 94.03 92.91 97.30 96.34 97.26 95.80 96.94 96.79 97.78
U. $\mathbb{C}_{u}$ 67.88/219.54 53.51/220.83 1.74/7.52 7.72/15.04 13.31/22.56 6.41/15.04 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 43.49/163.04 38.25/157.22 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 95.08 92.35 97.72 97.61 97.19 96.11 96.64 97.81 98.07
CIFAR-100 L. $\mathbb{C}_{u}$ 115.03/513.69 113.55/516.06 2.26/12.72 9.85/25.44 17.01/38.17 8.62/28.17 $\mathbb{Y}_{u}$ 91.90 91.50 98.00 97.70 98.00 97.60 97.60 98.00 98.10
$\mathbb{C}_{o}$ 68.54/285.63 60.44/287.24 1.08/5.28 4.43/10.56 7.56/15.84 4.19/14.32 $\mathbb{Y}_{o}$ 81.98 81.73 83.94 82.47 83.47 82.29 83.42 83.67 84.03
U. $\mathbb{C}_{u}$ 112.48/498.07 110.02/500.07 2.16/12.40 9.53/24.80 16.44/37.21 8.04/14.40 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 68.82/282.50 62.86/284.04 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 83.01 82.11 85.14 83.95 84.51 83.66 84.37 84.86 85.17
Hetero labels CIFAR-10 L. $\mathbb{C}_{u}$ 49.04/187.54 43.87/189.62 1.44/6.40 6.25/12.80 10.78/19.20 8.87/12.80 $\mathbb{Y}_{u}$ 93.20 93.90 99.40 99.80 99.80 99.70 99.50 99.90 99.90
$\mathbb{C}_{o}$ 50.31/187.53 44.34/189.63 1.46/6.40 6.28/12.80 10.78/19.21 7.37/12.82 $\mathbb{Y}_{o}$ 93.04 93.67 98.52 97.15 97.92 95.74 96.92 98.26 98.10
U. $\mathbb{C}_{u}$ 27.80/93.77 22.18/96.01 0.73/3.20 3.12/6.41 5.39/9.57 5.57/9.60 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 48.88/187.54 43.97/189.64 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 93.93 94.13 98.60 96.98 98.53 97.06 97.61 98.59 98.14
CIFAR-100 L. $\mathbb{C}_{u}$ 79.51/312.58 76.99/316.04 1.48/6.40 6.33/12.83 10.79/19.22 6.36/19.16 $\mathbb{Y}_{u}$ 90.30 92.10 99.90 99.70 98.90 98.70 99.00 99.00 99.10
$\mathbb{C}_{o}$ 79.74/312.60 77.02/316.03 1.50/6.42 6.35/12.80 10.83/19.22 5.78/11.21 $\mathbb{Y}_{o}$ 81.42 80.22 87.43 85.52 86.93 84.14 85.77 86.07 87.71
U. $\mathbb{C}_{u}$ 78.88/296.96 76.66/300.04 1.44/6.08 6.00/12.16 10.26/18.24 5.79/12.16 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 79.32/312.57 77.11/316.04 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 82.04 80.95 88.87 86.89 87.93 85.99 87.06 87.59 88.21
different tasks CIFAR-10 + MNIST L. $\mathbb{C}_{u}$ 93.61/381.11 88.92/384.05 2.95/12.80 13.50/25.61 31.64/37.29 13.08/24.76 $\mathbb{Y}_{u}$ 95.80 95.60 99.90 97.20 98.60 96.80 96.80 97.30 97.40
$\mathbb{C}_{o}$ 104.97/428.04 90.37/438.09 3.01/14.26 14.55/30.13 23.64/37.81 13.41/28.98 $\mathbb{Y}_{o}$ 95.13 95.73 98.64 96.71 97.15 95.09 96.47 96.31 96.71
U. $\mathbb{C}_{u}$ 80.97/286.96 61.15/288.04 2.39/9.60 9.89/19.20 14.20/26.13 9.92/19.08 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 103.79/428.09 90.35/437.94 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 95.12 95.74 99.07 96.89 97.01 95.33 96.34 96.56 96.72
$\varepsilon=0.1$ $\varepsilon=0.2$ $\varepsilon=0.1$ $\varepsilon=0.2$ $\varepsilon=0.1$ $\varepsilon=0.2$
  • Notation:Data distribution; ② Dataset; ③ Scenario (L. for learning, U. for unlearning); ④ Client type ($\mathbb{C}_{u}$ for clients involved in unlearning, $\mathbb{C}_{o}$ for clients excluded from unlearning); ⑤ Label type ($\mathbb{Y}_{u}$ for to-be-unlearned labels, $\mathbb{Y}_{o}$ for retained labels).
Table 3: Evaluation on Efficiency and Effectiveness over ResNet18 across Different SL Frameworks
(Client-side) Training / Transmission Time (s) Accuracy (%)
Existing SL SplitWiper SplitWiper+ Existing SL SplitWiper SplitWiper+
U-shaped Vanilla $\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$ U-shaped Vanilla $\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$
Clients with same task Homo labels Hetero volumes CIFAR-10 L. $\mathbb{C}_{u}$ 67.12/241.35 63.09/245.70 2.07/8.55 9.53/16.80 12.77/25.50 8.83/24.75 $\mathbb{Y}_{u}$ 90.90 88.60 89.70 89.20 90.30 88.30 90.40 89.90 91.20
$\mathbb{C}_{o}$ 42.06/159.30 38.69/177.45 1.25/5.55 6.29/11.10 8.15/16.65 5.51/9.60 $\mathbb{Y}_{o}$ 76.47 77.99 80.13 80.87 81.11 80.69 80.86 80.61 81.09
U. $\mathbb{C}_{u}$ 44.95/149.55 40.27/157.80 1.32/5.40 6.21/10.95 8.29/16.50 6.51/11.55 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 41.76/162.45 39.08/169.95 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 77.29 78.64 80.95 81.46 82.37 81.17 82.08 82.59 83.86
CIFAR-100 L. $\mathbb{C}_{u}$ 99.75/308.84 96.53/319.96 1.59/8.03 6.46/16.07 9.23/24.10 6.43/24.10 $\mathbb{Y}_{u}$ 89.30 88.90 92.25 91.70 93.20 92.00 93.10 93.00 94.60
$\mathbb{C}_{o}$ 79.82/300.96 77.22/307.48 1.53/6.08 6.03/12.13 8.80/18.21 6.05/6.08 $\mathbb{Y}_{o}$ 66.88 66.92 70.62 69.14 70.33 67.43 68.69 70.31 71.48
U. $\mathbb{C}_{u}$ 82.28/301.43 79.30/304.86 1.54/5.97 6.05/11.97 8.69/17.94 5.94/9.87 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 79.75/300.46 78.04/304.32 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 66.52 67.35 71.71 70.79 71.01 67.29 69.12 71.06 72.16
Partially overlapping labels CIFAR-10 L. $\mathbb{C}_{u}$ 79.72/268.65 75.56/295.35 2.43/10.05 11.66/20.25 17.80/30.30 13.69/28.05 $\mathbb{Y}_{u}$ 85.80 86.60 88.60 88.20 91.30 88.00 88.80 88.40 89.80
$\mathbb{C}_{o}$ 60.00/165.30 56.91/190.35 1.22/4.95 5.72/10.05 10.10/15.00 5.45/8.71 $\mathbb{Y}_{o}$ 77.15 78.56 80.86 80.44 80.84 80.26 80.53 80.61 81.62
U. $\mathbb{C}_{u}$ 55.97/178.20 53.18/207.45 1.73/7.05 8.02/14.10 12.50/21.15 8.75/18.60 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 60.34/169.80 52.24/198.45 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 80.02 80.67 81.69 81.49 82.51 81.24 82.11 81.81 82.45
CIFAR-100 L. $\mathbb{C}_{u}$ 134.15/497.29 129.23/499.94 2.62/12.40 11.18/24.80 17.18/37.25 10.80/36.89 $\mathbb{Y}_{u}$ 85.30 84.40 89.30 91.10 90.70 87.90 89.10 90.55 91.40
$\mathbb{C}_{o}$ 93.79/332.71 89.93/347.02 1.12/5.15 4.92/10.30 7.34/15.44 5.12/13.96 $\mathbb{Y}_{o}$ 63.88 63.03 69.25 68.55 69.09 68.02 68.86 69.29 70.29
U. $\mathbb{C}_{u}$ 126.87/452.21 120.65/459.62 2.42/10.96 10.16/23.52 15.58/35.41 10.16/14.98 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 95.28/336.73 89.05/346.28 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 64.52 63.81 70.46 69.01 69.72 68.58 68.73 70.33 71.03
Hetero labels CIFAR-10 L. $\mathbb{C}_{u}$ 46.88/176.40 43.52/208.05 1.46/6.15 7.19/12.30 9.31/18.15 7.89/15.90 $\mathbb{Y}_{u}$ 88.60 86.60 91.80 91.60 92.40 91.30 92.10 91.60 93.70
$\mathbb{C}_{o}$ 47.05/215.55 44.43/222.30 2.03/6.00 7.16/12.15 9.25/18.15 6.00/10.65 $\mathbb{Y}_{o}$ 85.96 85.66 86.23 87.38 88.53 86.92 87.66 88.02 88.37
U. $\mathbb{C}_{u}$ 24.96/88.20 22.63/104.10 0.72/3.00 3.63/6.00 4.61/9.00 4.03/7.80 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 47.11/205.95 44.40/215.25 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 86.42 87.22 87.92 87.94 88.72 87.12 88.14 88.17 89.04
CIFAR-100 L. $\mathbb{C}_{u}$ 83.54/312.00 79.94/321.64 1.46/6.40 5.72/12.80 9.29/19.20 6.67/19.20 $\mathbb{Y}_{u}$ 89.20 89.10 91.60 91.10 92.80 91.20 91.80 91.60 92.70
$\mathbb{C}_{o}$ 99.37/386.16 95.21/395.20 1.53/6.40 6.37/12.80 9.27/19.20 6.31/11.28 $\mathbb{Y}_{o}$ 69.61 70.68 73.76 71.96 72.56 70.99 71.09 72.60 72.93
U. $\mathbb{C}_{u}$ 74.51/295.36 71.32/300.12 1.41/6.08 6.00/12.16 8.89/18.24 6.10/12.16 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 98.15/386.20 94.85/393.00 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 69.99 70.57 74.72 72.27 73.29 71.98 72.38 72.59 73.33
different tasks CIFAR-10 + MNIST L. $\mathbb{C}_{u}$ 99.94/345.75 95.03/353.85 3.17/12.00 15.86/24.15 28.47/36.15 17.46/21.45 $\mathbb{Y}_{u}$ 91.30 90.40 94.10 92.50 92.60 91.60 92.30 93.00 93.10
$\mathbb{C}_{o}$ 108.48/391.80 100.52/397.95 3.13/13.35 16.53/26.85 21.69/40.20 17.75/28.95 $\mathbb{Y}_{o}$ 89.41 88.75 93.77 93.46 93.75 92.76 92.98 93.22 93.49
U. $\mathbb{C}_{u}$ 76.59/257.40 73.01/266.10 2.36/9.15 11.86/18.00 15.02/27.00 11.96/12.30 $\mathbb{Y}_{u}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$\mathbb{C}_{o}$ 107.99/392.55 98.81/397.95 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 $\mathbb{Y}_{o}$ 90.27 90.96 94.46 93.43 93.73 93.02 93.53 93.47 93.62
$\varepsilon=0.1$ $\varepsilon=0.2$ $\varepsilon=0.1$ $\varepsilon=0.2$ $\varepsilon=0.1$ $\varepsilon=0.2$
  • Notation:Data distribution; ② Dataset; ③ Scenario (L. for learning, U. for unlearning); ④ Client type ($\mathbb{C}_{u}$ for clients involved in unlearning, $\mathbb{C}_{o}$ for clients excluded from unlearning); ⑤ Label type ($\mathbb{Y}_{u}$ for to-be-unlearned labels, $\mathbb{Y}_{o}$ for retained labels).
Model architecture. We utilize the architecture based on a simplified VGG [28] and a ResNet18 [29] to construct the client-side and server-side models for our split (un)learning experiments. In this configuration, the first convolutional layer of VGG or ResNet18 serves as the client model $\mathcal{F}_{o}^{k}$, while the remaining other layers comprise the server model $\mathcal{F}_{o}^{s}$. Given that VGG and ResNet have negligible differences in their first convolutional layer, we let the number of parameters in the client model $\mathcal{F}_{o}^{k}$ be consistent across both architectures. The number of epochs for client-side pre-training is set to $N=10$. The number of server-side model training and retraining epochs is set to $M=30$ for experiments involving CIFAR-10 and $M=50$ for experiments utilizing CIFAR-100.
In scenarios where clients train on the same task, each client’s model, $\mathcal{F}_{o}^{k},\forall k$, adheres to a uniform structure with the same input and output sizes across all clients. Conversely, in scenarios with clients engaged in different tasks, the input size of each $\mathcal{F}_{o}^{k}$ is tailored based on the specific training dataset $D_{o}^{k}$ used locally. However, the output size remains consistent to meet the input requirements of the server model $\mathcal{F}_{o}^{s}$. This design ensures that despite the diverse data types and tasks, all client models can interface correctly with the centralized server model within the SL framework.
Benchmarks and experimental groups. We employ existing Vanilla and U-shaped SL [4] as benchmarks to evaluate the performance of SplitWiper. To ensure a fair comparison, we initialize Vanilla and U-shaped SL with the pre-trained results from SplitWiper. Then, both the client and server in these existing frameworks are set to undergo the same $M$ epochs as specified in SplitWiper to train in the learning task or re-train when asking for unlearning. By aligning the initial conditions and training durations across all frameworks, we assess the efficiency and effectiveness of SplitWiper relative to the established benchmarks.
To assess the impact of different label expansion strategies in SplitWiper and SplitWiper+, we categorize the experiments based on various expansion factors. Specifically, four groups are created: three with uniform label expansions of $\gamma=1$, $2$, or $3$. When $\gamma=1$ (SplitWiper), labels are shuffled and re-indexed, but their actual count remains unmasked, similar to Vanilla SL where real labels are shared. For $\gamma=2$ or $\gamma=3$ (SplitWiper+), clients expand labels to $\gamma$ instances and apply $\varepsilon$-DP to protect the corresponding intermediate values, with $\varepsilon=0.1$ and $0.2$. We also experiment with a scenario denoted as $\gamma=\Gamma_{p}$ where $\gamma_{q}$ for each $Y_{q}$ is randomly selected within a given range $\Gamma_{p}=[1,3]$. This method allows us to examine how variability in the expansion factor affects label privacy and system performance.
Performance metrics. We use training and transmission times as primary indicators of computational and communication overheads for different SL frameworks to highlight the superiority of the proposed one-way-one-off propagation. We focus on client-side overhead due to the comparable server-side computational costs across frameworks. In SplitWiper and SplitWiper+, the server-side complexity is $O(M\cdot\mathcal{H}_{S})$, where $M$ is the number of training epochs and $\mathcal{H}_{S}$ is the complexity of a single server training iteration. Similarly, Vanilla SL has a server-side complexity of $O(M\cdot\mathcal{H}_{S})$, while U-shaped SL reduces this to $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$ due to the output layer’s location in the client model. We consider a bandwidth of $100$ Mbps to simulate transmission costs between clients and the server. We evaluate the unlearning effectiveness of each framework by comparing the label accuracy during both the learning and unlearning phases. Our experiments run on NVIDIA PCIe A100, 2 $\times$ 40GB, and we use Python 3.7.16 and PyTorch 2.3.1 to build and train the proposed framework.
5.2 Unlearning Efficiency and Overhead
((a)) Training time over VGG
((b)) Transmission time over VGG
((c)) Training time over ResNet18
((d)) Transmission time over ResNet18
Figure 3: Client-side time consumption across different SL frameworks.
((a)) Training and transmission times across varies $\gamma$ over VGG
((b)) Training and transmission times across varies $\gamma$ over ResNet18
((c)) Test accuracy across varies $\varepsilon$ under different $\gamma$ settings over VGG
((d)) Test accuracy across varies $\varepsilon$ under different $\gamma$ settings over ResNet18
Figure 4: Impacts of the expansion factor $\gamma$ and privacy budget $\varepsilon$ on SplitWiper and SplitWiper+ in terms of client-side training and transmission time consumption and test accuracy.
The left sides of Table 5 and Table 3 show the training and transmission times for clients in various SL frameworks over VGG and ResNet18, focusing on clients with unlearning requests ($\mathbb{C}_{u}$) and other clients ($\mathbb{C}_{o}$) across both learning and unlearning scenarios. Since the degree of DP-based noise addition does not affect time costs, we exclude the influence of $\varepsilon$ from our analysis, focusing solely on the effects of different label expansion factors $\gamma$.
The results highlight the efficacy of SplitWiper in achieving G1 (independent unlearning). Training and transmission times in SplitWiper are significantly lower than those in Vanilla and U-shaped SL across all data distributions and scenarios. Crucially, during unlearning, only the requesting clients, $\mathbb{C}_{u}$, incur additional costs, while other clients, $\mathbb{C}_{o}$, remain unaffected, as the server uses cached data from the learning phase to handle unlearning, minimizing overall client-side expenses and disruptions. The time overhead for the label expansion strategy discussion between clients (cf. Algorithm 3) is negligible, measured in milliseconds, and does not significantly impact our analysis of training and transmission times.
In SplitWiper and SplitWiper+, clients only need to train the pre-trained model over local data once. They then mask labels according to the label expansion strategy and apply DP protection to the corresponding intermediate values before transmitting them to the server in a one-way propagation manner. Subsequent training phases focus exclusively on the server, disconnecting from client-side computations. Therefore, the training and transmission times for clients are fixed values dependent on $\gamma$ and remain unaffected by the number of training epochs $M$. In contrast, in Vanilla and U-shaped SL, where each epoch involves client participation, the computational and transmission times increase linearly with $M$, as listed in Table 1 (see Appendix B).
Fig. 3 shows the average training and transmission times for all clients across different SL frameworks as training epochs $M$ increase. At $M=1$, the times for Vanilla and U-shaped SL are nearly identical to those of SplitWiper with $\gamma=1$. As $\gamma$ increases (SplitWiper+), computational and communication costs rise, but they typically stay below $5$ due to client-specified balancing. Even in practical settings with $M=30$ for CIFAR-10 and CIFAR-10/MNIST, and $M=50$ for CIFAR-100, SplitWiper retains a clear advantage, especially during unlearning, where only the requesting clients ($\mathbb{C}_{u}$) incur costs, while others ($\mathbb{C}_{o}$) do not.
To intuitively demonstrate the impact of the label expansion factor $\gamma$ on client-side overhead in SplitWiper+, we plot the average training and transmission times for all clients across different $\gamma$ values in the learning scenarios in Fig. 4(a) and Fig. 4(b). Both figures, whether over VGG or ResNet18, show a roughly linear increase in overhead as $\gamma$ increases. The case where $\gamma=\Gamma_{p}$ represents the scenario where $\gamma_{q},\forall q$ is randomly selected as an integer from the range $[1,3]$, with the average $\gamma_{q}$ being $1.9$ across our experiments. This scenario incurs slightly higher overhead due to the variability in $\gamma_{q}$ compared to a uniform expansion factor. Since CIFAR-100 and CIFAR-10 are similar in size ($\sim 177$ MB), their computational and transmission overheads are nearly identical. However, using CIFAR-10 and MNIST in scenarios with different client tasks significantly increases training and transmission times.
Takeaways. SplitWiper/SplitWiper+ eliminate overhead increases as epochs grow through one-way-one-off propagation, ensuring that unrelated clients remain unaffected during unlearning. Despite label expansion overhead, SplitWiper+ reduces computational and communication costs by as much as SplitWiper, achieving over 97-99% reduction in unlearning compared to existing SL methods.
5.3 Unlearning Effectiveness and Accuracy
The right sides of Table 5 and Table 3 present the accuracy of the label to be unlearned, $Y_{u}$, and other labels, $Y_{o}$, under both learning and unlearning scenarios across the test set. It is evident that all frameworks, including existing ones and ours with varies $\gamma$ and $\varepsilon$, effectively meet G2 (effective unlearning with retained utility) by reducing the accuracy of $Y_{u}$ to zero through unlearning, while simultaneously ensuring that the accuracy of $Y_{o}$ remains unaffected. Considering the CIFAR-100 dataset features a larger number of labels and fewer training examples per label, it poses a higher learning challenge compared to CIFAR-10, resulting in generally lower accuracy across all frameworks on CIFAR-100. Additionally, accuracy on CIFAR-100 is lower with the ResNet18 architecture compared to VGG, as VGG demonstrates better suitability for this specific task in our experiments.
((a)) Perfect clustering accuracy with elbow method.
((b)) Inclusive clustering accuracy with elbow method.
((c)) Perfect clustering accuracy with real labels number $Q$.
((d)) Inclusive clustering accuracy with with real labels number $Q$.
Figure 5: KMeans inclusive clustering accuracy and perfect clustering accuracy with Variable $\gamma$ and $\varepsilon$, using the elbow method or the real label number $Q$ as the cluster parameter. The low perfect clustering accuracy with both the elbow method and using $Q$ as the parameter indicates the effective preservation of label-related privacy in SplitWiper+.
Compared to Vanilla and U-shaped SL, SplitWiper shows a notable accuracy increase for both $Y_{u}$ and $Y_{o}$ in the learning scenario. It surpasses Vanilla SL by 4.77% in CIFAR-10, 4.37% in CIFAR-100, and 2.98% in scenarios involving CIFAR-10 and MNIST. Against U-shaped SL, SplitWiper achieves accuracy improvements of 4.18% in CIFAR-10, 3.82% in CIFAR-100, and 3.54% in mixed CIFAR-10 and MNIST scenarios. The key feature of SplitWiper is freezing client parameters after pre-training, keeping them unaffected during server-side training. When all clients share a server model $\mathcal{F}_{o}^{s}$, back-propagation from $\mathcal{F}_{o}^{s}$ can cause overfitting or hinder learning in client-$k$’s model $\mathcal{F}_{o}^{k}$. Freezing client parameters in SplitWiper enhances efficiency, reduces overhead, and minimizes cross-client interference, improving the accuracy of $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$.
Comparing SplitWiper and SplitWiper+, Figs. 4(c) and 4(d) show the impact of $\varepsilon$ and $\gamma$ on accuracy. As $\gamma$ increases, accuracy decreases slightly but remains within acceptable limits. In SplitWiper+ with $\gamma>1$, higher $\varepsilon$ reduces noise and improves test accuracy. SplitWiper+ with $\varepsilon=0.2$ can even outperform SplitWiper ($\gamma=1$) due to minor noise and label expansion acting as data augmentation [30], enhancing accuracy and generalization. Specifically, SplitWiper+ with $\varepsilon=0.2$ and $\gamma=\Gamma_{p}$ shows an average accuracy increase of 0.37% in CIFAR-10 and 0.12% in CIFAR-100, particularly for $\mathbb{Y}_{o}$, while maintaining comparable performance on $\mathbb{Y}_{u}$. This demonstrates SplitWiper+’s ability to enhance privacy protection without compromising utility.
Takeaways. SplitWiper and SplitWiper+ effectively unlearn upon request while preserving the accuracy of retained data. The unlearning process in SplitWiper is over 108% more effective than existing SLs, with SplitWiper+ further enhancing accuracy by an additional 0.37%.
5.4 Label-related Privacy Analysis
Considering G3 (privacy-preserving label sharing), we evaluate the server’s ability to infer real label characteristics from the received masked labels and corresponding intermediate outputs. Given that the server lacks access to a local dataset and the shared labels are anonymized, it faces significant constraints in conducting semantic inference attacks on labels. We focus our analysis on the server’s capability to deduce the actual number of labels used by clients, which involves cracking the critical mapping mechanism $(\mathcal{G}_{Y},\mathcal{G}_{V}):\{Y_{q},V_{q}^{k}\}\rightarrow\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$. This mapping is the key differentiator between SplitWiper and SplitWiper+, highlighting the enhanced privacy features of the former in protecting label information against server-based inference attacks.
In SplitWiper+, the server employs clustering algorithms on the received data pairs $\{Y_{x}^{*},{V_{x}^{*}}^{k}\}$ to group the labels with closer corresponding intermediate outputs into the same cluster, hypothesizing that these labels are expanded from the same original label. We assess the effectiveness of this clustering-based label matching attack using two metrics:
  • $\bullet$
    Inclusive clustering accuracy. This accuracy indicates the probability that labels derived from the same real label are inclusively grouped within the same cluster formed by the clustering algorithm. It underscores the containment relationship where a single cluster may encompass an entire real label group, but not necessarily align exactly with it.
  • $\bullet$
    Perfect clustering accuracy. This accuracy indicates the proportion of instances where the clustered groups formed by the clustering algorithm match exactly with the true label expansion groups, reflecting a flawless alignment.
Inclusive clustering accuracy assesses whether the distances among masked labels, derived from the same original label, are close enough to be accurately grouped by the clustering algorithm. This metric checks if related labels cluster together based on their proximity in feature space. In contrast, perfect clustering accuracy also considers whether there is sufficient separation between different groups, evaluating the algorithm’s ability to clearly distinguish between them. Together, these metrics offer insights into SplitWiper+’s effectiveness in preserving label-related privacy by maintaining the integrity of masked label structures during clustering.
We employ the KMeans clustering algorithm [31], a standard method renowned for its robustness and widespread use in detecting patterns in ungrouped data. Given that the real number of labels, $Q$, is confidential to the server, we utilize the elbow method [31] to determine an appropriate number of clusters $K$. Additionally, to illustrate the potential privacy breach in SplitWiper+, we also conduct experiments where $Q$ is used directly as the parameter $K$ for KMeans. This scenario simulates the impact of a hypothetical leakage of $Q$, revealing how such exposure could compromise the privacy measures implemented within SplitWiper+.
Given that the pre-trained models on each client are consistent across both VGG and ResNet18 architectures—since resource-limited clients have few layers and training techniques like dropouts and batch normalization are applied on the server side—the intermediate values show no significant variation. Therefore, we do not differentiate between VGG and ResNet18 in our experiments, focusing instead on the impact of intermediate values on label privacy. As shown in Fig. 5(a), when $Q$ is unknown and the server estimates the optimal cluster number using the elbow method, perfect clustering accuracy remains low, not exceeding 10% with $\gamma=3$. This suggests that SplitWiper+, using a label expansion strategy with $\gamma=3$, preserves over 90% of label-related privacy, effectively safeguarding sensitive information against potential inference attacks. As $\gamma$ increases, or with non-uniform expansion factors ($\gamma=\Gamma_{p}$), perfect clustering accuracy decreases, further enhancing privacy protection.
In contrast, the inclusive clustering accuracy (Fig. 5(b)), is notably higher than the perfect clustering accuracy, due to its less stringent criteria. As more masked labels are clustered together by KMeans, the likelihood of satisfying the inclusive clustering accuracy increases. Therefore, under conditions of higher $\gamma$ or in SplitWiper with $\gamma=1$, when more masked labels are grouped into the same cluster, the inclusive clustering accuracy rises. However, this scenario primarily indicates that masked labels derived from the same original label are grouped together, without providing the means to differentiate between distinct groups of labels. While this may reveal minimal related information, it is insufficient for practical use.
When the number of real labels $Q$ is disclosed to the server (Fig. 5(c)), clustering-based label matching attacks achieve higher perfect clustering accuracy, increasing by an average of 16% compared to when $Q$ remains confidential. This increase highlights the additional privacy risks posed by the exposure of label counts, underscoring the necessity of implementing privacy-preserving measures in SplitWiper+. Furthermore, a larger $\gamma$ exacerbates the risk related to the leakage of $Q$, as a higher expansion factor results in a greater number of masked labels derived from each original label, providing more information to assist clustering when $Q$ is known.
Takeaways. SplitWiper+ protects label privacy and corresponding intermediate values through expansion and DP-based obfuscation, preventing server inference attacks via clustering. When the true label count is concealed, SplitWiper+ preserves over 90% of label privacy, and even when exposed, it still offers around 70% protection.
6 Related Work
Unlearning is challenging in distributed environments. Machine unlearning has not been fully explored within current distributed learning frameworks, including SL. While efficient unlearning strategies have gained attention in FL, which also involves collaboration between clients and a central server. Some researchers have worked on accelerating retraining across all clients, the most robust unlearning method in FL, using optimized algorithms like distributed Newton-type updates [32], SGA-EWC [33], and SFU [34]. Solutions like FedEraser [35] reduce the impact on non-involved clients through calibration but still require cooperation from some clients. However, in SL, where client resources are often limited, non-requesting clients may lack the incentive or ability to contribute, especially if they are uninterested or unable to participate due to network issues. This presents a significant challenge in SL, where developing unlearning methods that accommodate these constraints remains unresolved.
Existing SISA is inadequate to SL scenarios. SISA [6] is a state-of-the-art solution for robust and efficient machine unlearning, enabling rapid adaptation to data removal requests without requiring full model retraining. By strategically partitioning data into distinct “shards” linked to different model versions, SISA allows for the targeted unlearning of specific data shards without requiring the participation of all shards. This method has been successfully applied in various domains, including unlearning in graph neural networks [13] and large language models [36]. However, SISA-based unlearning faces significant challenges in multi-client SL scenarios due to the interdependence between client model segments and the shared server model segment, which violates the Isolated requirement of SISA. Consequently, the direct application of conventional SISA is impractical for existing SL scenarios.
Label privacy is important in SL unlearning. Beyond unlearning, privacy concerns regarding intermediate outputs and labels in SL have been widely studied. U-shaped SL [4] avoids label propagation to the server by placing the output layer on the client side but violates SISA’s Isolated requirement by necessitating the return of intermediate values across all clients, disabling the independent unlearning and adding overhead to resource-limited clients. If labels are shared but privacy is required, existing research primarily addresses privacy risks from transmitting intermediate outputs to the server. Wu et al.[37] introduced a DP mechanism to perturb these outputs, while Li et al.[20] developed the Marvell method to quantify privacy leakage and optimize noise structures. However, these methods overlook the exposure of sensitive information through label quantity and semantics. Xiao et al. [38] obfuscates both labels and outputs via multiple linear combinations, reducing association with original data but still fails to conceal label quantity, potentially exposing sensitive information.
Ours. Our design overcomes the mentioned challenges. SplitWiper adapts SISA for multi-client SL, starting with low-effort pre-training on the clients followed by freezing their weights. This allows for one-way-one-off propagation to the server, enabling efficient unlearning without involving non-participating clients. Additionally, SplitWiper+ enhances label privacy by expanding, shuffling, and anonymizing real labels, with intermediate outputs expanded and protected using a DP-based mechanism.
7 Conclusion
In this work, we achieve the first practical split unlearning framework. We presented SplitWiper and SplitWiper+, which designed an one-way-one-off propagation to decouple propagation between clients and between clients and the server, enabling SISA-based unlearning with absent clients. SplitWiper+ further enhanced client label privacy against the server by employing DP. Experiments across diverse data distributions and tasks demonstrated that SplitWiper effectively achieved unlearning by involving only the requesting clients, resulting in 0% unlearning accuracy and over 108% greater effectiveness on retained data compared to existing SL methods. Additionally, it maintained constant overhead and reduced computational and communication costs by over 99%. SplitWiper+ preserved over 90% of label privacy, effectively protecting sensitive information from the server.
Appendix
Appendix A Proof of Theorem 1
Proof (Theorem 1).
$\forall{V_{a}^{*}},{V_{b}^{*}}\in\mathbb{V}_{exp}$ and $\forall{V}_{t}\in\mathbb{V}$,
$e^{-\varepsilon}\leq\frac{\Pr[{V}_{t}|{V_{a}^{*}}]}{\Pr[{V}_{t}|{V_{b}^{*}}]}\leq e^{\varepsilon}.$ (3)
Thus, the probability that any two elements in $\mathbb{V}_{exp}$ correspond to the same element in $\mathbb{V}$ can be evaluated. Use $\boxtimes$ to indicate that two masked elements correspond to the exact origin one. $\forall{V_{a}^{*}},{V_{b}^{*}},{V_{c}^{*}}\in\mathbb{V}_{exp}$, according to (3), we have
$\frac{\Pr[{V_{a}^{*}}\boxtimes{V_{b}^{*}}]}{\Pr[{V_{a}^{*}}\boxtimes{V_{c}^{*}}]}$ $=\frac{\sum_{q=1}^{Q}{\Pr[{V}_{t}|{V_{a}^{*}}]\Pr[{V}_{t}|{V_{b}^{*}}]}}{\sum_{q=1}^{Q}{\Pr[{V}_{t}|{V_{a}^{*}}]\Pr[{V}_{t}|{V_{c}^{*}}]}}$ (4)
$\leq\frac{\sum_{q=1}^{Q}{\Pr[{V}_{t}|{V_{a}^{*}}]((e^{\varepsilon}\Pr[{V}_{t}|{V_{a}^{*}}])}}{\sum_{q=1}^{Q}{\Pr[{V}_{t}|{V_{a}^{*}}]((e^{-\varepsilon}\Pr[{V}_{t}|{V_{a}^{*}}])}}$
$=e^{2\varepsilon}.$
Thus, the probability that any two elements in $\mathbb{V}_{exp}$ correspond to the same element in $\mathbb{V}$ satisfies $2\varepsilon$-DP.
As demonstrated in Theorem 1, the server cannot infer the mapping $\mathcal{G}_{V}$ from the differences between any two DP-protected intermediate values. This inability also prevents the server from deducing the mapping $\mathcal{G}_{Y}$ from the intermediate values, thereby protecting the original label information from being derived from the masked labels. Within the SplitWiper framework, carefully selecting the parameter $\varepsilon$ is essential to balance privacy with maintaining data utility and model effectiveness.
Table 4: Notation
(Un)Learning client-$k$ The $k$-th client
$D_{o}^{k}$ Original dataset owned by client-$k$
$D_{u}^{k}$ Updated dataset (client-$k$) for unlearning
$(X_{i},Y_{i})$ The $i$-th data sample and its corresponding label
$\mathcal{F}_{o}^{k}/\mathcal{F}_{o}^{s}$ Original model maintained by client-$k$ / the server
$\mathcal{F}_{u}^{k}/\mathcal{F}_{u}^{s}$ Updated model (client-$k$ / server) for unlearning
$W_{k}^{(t)}$ Model weights of client-$k$ at iteration $t$
$W_{s}^{(t)}$ Model weights of the server at iteration $t$
$\mathcal{L}_{k}$ Loss function of client-$k$
$\mathcal{T}$ Evaluation function that assesses models
$N$ Average epochs of pre-training for clients in training/retraining
$M$ Average epochs of training for server in training/retraining
$\mathbb{C}_{u}$ Set of clients involved during unlearning
$\mathbb{C}_{o}$ Set of clients not involved during unlearning
Label Expansion $Q$ Number of real labels
$\mathbb{Y}$ Complete set of real labels
$\mathbb{Y}_{k}$ Set of labels owned by client-$k$
$\mathbb{Y}_{exp}$ Set of expanded labels
$\mathcal{G}_{Y}$ Corresponding map between $\mathbb{Y}$ and $\mathbb{Y}_{exp}$
$\gamma_{q}$ Label expansion factor for the $q$-th label, $Y_{q}$
$V_{q}^{k}$ Intermediate values of $Y_{q}$ on client-$k$
$\varepsilon$ Privacy budget of the DP mechanism
Algorithm 1 Learning in SplitWiper
1:$K$ (Number of clients), $N$ (number of epochs for client model training), $M$ (number of epochs for server model training);
2:Predicted label
3:Clients: (Conducted in parallel for each client-$k$)
4:for $k=1$ to $K$ do
5:    Initialize client model $\mathcal{F}_{o}^{k}$
6:    for $n=1$ to $N$ do
7:        Train $\mathcal{F}_{o}^{k}$ on $D_{o}^{k}$, excluding server
8:    end for
9:    Freeze model weights of $\mathcal{F}_{o}^{k}$, disabling gradient updating functions
10:    Send labels and intermediate outputs to the server
11:end for
12:Server:
13:Cache received labels and intermediate outputs
14:Initialize server model $\mathcal{F}_{o}^{s}$
15:for $m=1$ to $M$ do
16:    Train $\mathcal{F}_{o}^{s}$ on cached labels and intermediate outputs
17:end for
18:RETURN predicted label
Algorithm 2 Unlearning in SplitWiper
1:$N$ (number of epochs for client model retraining), $M$ (number of epochs for server model updating), initiating client-$k$
2:Predicted label
3:Clients: (Only client-$k$ involved, others remain silent)
4:Unfreeze gradient updating functions
5:Initialize local model $\mathcal{F}_{u}^{k}$
6:Remove intended unlearned samples from $D_{o}^{k}$, obtaining $D_{u}^{k}$
7:for $n=1$ to $N$ do
8:    Train $\mathcal{F}_{u}^{k}$ on $D_{u}^{k}$, excluding server
9:end for
10:Freeze updated weights of $\mathcal{F}_{u}^{k}$
11:Send intermediate outputs of $\mathcal{F}_{u}^{k}$ and corresponding labels to the server
12:Server:
13:Update cache with intermediate outputs of $\mathcal{F}_{u}^{k}$
14:Initialize server model $\mathcal{F}_{o}^{s}$ based on cache, $\mathcal{F}_{u}^{s}\leftarrow\mathcal{F}_{o}^{s}$
15:for $m=1$ to $M$ do
16:    Train $\mathcal{F}_{u}^{s}$ on cached records from client-$k$ and other clients
17:end for
18:RETURN predicted label
Algorithm 3 Label Expansion Strategy
1:$K$ (number of clients), $\{\mathbb{Y}_{1},\cdots,\mathbb{Y}_{K}\}$ ( clients’ local label sets), $\mu$ ( maximum number of shared labels in each round ), $\lambda_{1}/\lambda_{2}$ ( threshold of client participation ), $\Gamma$ (selected range for label expansion factor).
2:Real label set $\mathbb{Y}$, Expanded label set $\mathbb{Y}_{exp}$, and corresponding map between these two set $\mathcal{G}_{Y}$
3:Clients:
4:Initialize two empty sets $\mathbb{Y}$ and $\mathbb{Y}_{exp}$, an empty mapping $\mathcal{G}_{Y}$, Initialize array $\mathbb{J}$ with $K$ zeros
5:repeat
6:    for $k=1$ to $K$ do
7:        if $\mathbb{J}[k]\neq 0$ then continue
8:        if generate a random number $r_{1}<\lambda_{1}$ then
9:              Select $r_{\mu}\leftarrow[1,\min(\max(|\mathbb{Y}_{k}\setminus\mathbb{Y}|,1),\mu)]$ elements from $\mathbb{Y}_{k}\setminus\mathbb{Y}$ and update $\mathbb{Y}\leftarrow\mathbb{Y}\bigcup$ the selected elements
10:        if generate a random number $r_{2}<\lambda_{2}$ and $\mathbb{Y}_{j}\subseteq\mathbb{Y}$ then $\mathbb{J}[k]\leftarrow 1$
11:    end for
12:until $\sum_{k=1}^{K}\mathbb{J}[k]=K$
13:for $q=1$ to $\left|\mathbb{Y}\right|$: Generate $\gamma_{q}$ from $\Gamma$ and add $\mathbb{Y}[q]$ for $\gamma_{q}$ times to $\mathbb{Y}_{exp}$
14:$\mathbb{Y}_{exp}\leftarrow$ Shuffle the elements in $\mathbb{Y}_{exp}$
15:for $q=1$ to $\left|\mathbb{Y}\right|$: $\mathcal{G}_{Y}[\mathbb{Y}[q]]\leftarrow$ List($\mathbb{Y}_{exp}.\text{index}(\mathbb{Y}[q])$)
16:$\mathbb{Y}_{exp}\leftarrow$ Re-number the elements in $\mathbb{Y}_{exp}$ according to the index
17:RETURN $\mathbb{Y}$, $\mathbb{Y}_{exp}$, $\mathcal{G}_{Y}$
Algorithm 4 DP-based Expansion-and-Masking Scheme
1:$\mathcal{G}_{Y}$ (map between $\mathbb{Y}$ and $\mathbb{Y}_{exp}$), $\varepsilon$ (privacy budget of DP), $\{Y_{q},V_{q}^{k}\}$ (pairs of labels and intermediate values from client-$k$).
2:Output set of the expansion-and-masking scheme $\mathbb{U}_{exp}=\{\mathbb{Y}_{exp},\mathbb{V}_{exp}\}=\bigcup\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$
3:Clients:
4:Initialize an empty set $\mathbb{U}_{exp}$
5:$X\leftarrow\mathcal{G}_{Y}[Y_{s}]$
6:for $x\in X$  do
7:    $Y_{e}^{*}\leftarrow x$
8:    ${V_{e}^{*}}^{k}\leftarrow V_{q}^{k}+DP(\varepsilon)$
9:    $\mathbb{U}_{exp}\leftarrow\mathbb{U}_{exp}\bigcup\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$
10:end for
11:RETURN $\mathbb{U}_{exp}$
Appendix B Complexity Analysis
To further demonstrate the advantages of the SplitWiper and SplitWiper+ frameworks in reducing computational and communication overhead, we analyze its complexity in both learning and unlearning scenarios. We compare these metrics with traditional Vanilla and U-shaped SL, which adhere to the standard procedures of the usual round-robin SL processes detailed in [4], as shown in Table 1.
B.1 Learning Scenarios
B.1.1 Computational Complexity
Client-side computational complexity. In the SISA-based SplitWiper framework, the primary computational cost is derived from the pre-training of local models on each client-$k$, which is $O(N\cdot\mathcal{H}_{K})$, where $N$ is the epochs for client model pre-training and $\mathcal{H}_{K}$ is the complexity of a single training iteration. While in SplitWiper+, beyond the pre-training computational cost, there is an additional component that arises from the computations required by the label expansion strategy implemented after the pre-training phase on the client side. This additional cost is $O(\gamma\;\mathcal{H}_{exp})$, where $\mathcal{H}_{exp}$ represents the complexity of the label expansion-and-masking scheme, and $\gamma$ is the average of the label expansion factors.
However, since the pre-training of client models is conducted independently of the server, it is excluded from our computational complexity analysis for the SL process. Instead, we focus on the final epoch of the pre-training process, disregarding the impact of $N$. Due to the one-way-one-off propagation design, clients freeze their local models after pre-training and do not participate in the subsequent $M$ epochs of server training. Therefore, the computational complexity of the proposed frameworks is also unaffected by $M$. For a total of $K$ clients, the computational complexity across all clients in SplitWiper is quantified as $O(K\cdot\mathcal{H}_{K})$. Correspondingly, the computational complexity of SplitWiper+ is $O(K\cdot(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp}))$.
In contrast, traditional SL, such as Vanilla and U-shaped SL, requires all $K$ clients to be involved across the entire learning process for $M$ epochs, leading to higher computational complexity. Specifically, the computational complexity for the Vanilla SL framework is $O(M\cdot K\cdot\mathcal{H}_{K})$. As for the U-shaped framework, which allocates the output layer to the client side, an additional computation cost is required. Consequently, the total complexity for the U-shaped framework is $O(M\cdot K\cdot(\mathcal{H}_{K}+\mathcal{H}_{out}))$, with $\mathcal{H}_{out}$ representing the complexity associated with the output layer operations.
In Algorithm 4, the complexity of the expansion-and-masking scheme $\mathcal{H}_{exp}$ scales linearly with the number of real labels and the size of intermediate values. This complexity remains within the same order of magnitude as $\mathcal{H}_{K}$, especially given the significant size of large models today. This computational load drives clients to use SL in collaboration with a server rather than independently handling the entire learning process. For security and efficiency, the expansion factor $\gamma$ is typically kept small, and training epochs $M$ are much larger, often ranging from several dozen to hundreds. As a result, the computational overhead of the SplitWiper+ framework is substantially lower than that of traditional Vanilla and U-shaped SL frameworks. Moreover, SplitWiper has only $1/M$ of the computational complexity of Vanilla SL, further highlighting its advantages.
Server-side computational complexity. In SplitWiper and SplitWiper+, the server’s updating of its model incurs a computational cost of $O(M\cdot\mathcal{H}_{S})$, where $M$ is the number of epochs for server model training and $\mathcal{H}_{S}$ is the complexity of a single server training iteration. This complexity mirrors that in the Vanilla SL framework, where the server-side complexity is also $O(M\cdot\mathcal{H}_{s})$. Conversely, the U-shaped SL framework generally exhibits slightly lower server-side complexity, which is $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$, because clients undertake output layer operations, thus offloading some of the computational demand from the server.
B.1.2 Communication Overhead
Client-side communication complexity. In the SplitWiper and the SplitWiper+ frameworks, the communication overhead is determined by transmitting the (masked) labels and corresponding intermediate outputs from each client to the server in a one-way propagation manner. This leads to a complexity of $O(K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ for SplitWiper and $O(K\cdot\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ for SplitWiper+, where $\mathcal{I}_{K}$ is the size of intermediate outputs and $\mathcal{I}_{Y}$ is the size of labels on client-side.
Compared to Vanilla SL, which incurs a per-epoch communication complexity of $O(K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ and U-shaped SL that does not require label transmission, our label expansion strategy in SplitWiper+ does introduce additional communication overhead. However, considering that traditional SL frameworks involve continuous communication between clients and the server across $M$ epochs, the overall communication complexities are respectively $O(M\cdot K\cdot(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ and $O(M\cdot K\cdot\mathcal{I}_{K})$. Given that $M$ is significantly greater than $\gamma$, the SplitWiper+ framework still maintains a substantial advantage regarding communication efficiency.
((a)) CIFAR-10 over VGG
((b)) CIFAR-100 over VGG
((c)) CIFAR-10 + MNIST over VGG
((d)) CIFAR-10 over ResNet18
((e)) CIFAR-100 over ResNet18
((f)) CIFAR-10 + MNIST over ResNet18
Figure 6: Test accuracy vs epochs for different SL frameworks over the VGG and ResNet18 architectures across various datasets with heterogeneous label distributions. As the number of training epochs ($M$) increases, the accuracy of SplitWiper and SplitWiper+ converges.
Server-side communication complexity. In the proposed frameworks, the information transfer from clients to the server follows a one-way propagation manner, allowing us to consider that the server incurs no communication costs. In contrast, the Vanilla SL framework requires the server to send the loss back to each client to facilitate model training, leading to a communication complexity of $O(M\cdot\mathcal{I}_{loss})$ over $M$ epochs, where $\mathcal{I}_{loss}$ represents the size of the loss during each epoch. Similarly, the U-shaped SL framework necessitates the server to return outputs back to clients for operations related to the final output layer, resulting in a communication complexity of $O(M\cdot\mathcal{I}_{S})$, where $\mathcal{I}_{S}$ is the size of the server-side outputs.
B.2 Unlearning Scenarios
In the unlearning scenario, the retraining process in traditional Vanilla and U-shaped SL frameworks necessitates the participation of all clients. This implies that each client must re-initiate model updates and transmit intermediate values for an additional $M$ epochs, replicating the computational and communication costs observed in the learning scenario.
Conversely, SplitWiper and SplitWiper+ introduce an optimized approach, wherein unlearning requests are isolated to the specific client initiating the request, significantly reducing the overall burden. This isolation avoids the need for unrelated clients to engage in redundant computations and communications, thereby enhancing efficiency and reducing client-side total resource consumption. In terms of client-side computational complexity in SplitWiper framework, the primary costs come from the retraining of the local model on the initiating client-$k$, which is $O(\mathcal{H}_{K})$. In contrast, in SplitWiper+, client-$k$ also incurs the additional cost of performing a new round of the expansion-and-masking scheme, resulting in a computational complexity of $O(\mathcal{H}_{K}+\gamma\;\mathcal{H}_{exp})$.
The communication overhead is determined by transmitting the intermediate outputs from the initiating client-$k$ to the server, leading to a complexity of $O(\gamma\;(\mathcal{I}_{K}+\mathcal{I}_{Y}))$ in the proposed frameworks, where $\gamma=1$ represents SplitWiper, and $\gamma>1$ represents SplitWiper+. It satisfies G1 (independent unlearning), thus improving the robustness and reducing the overhead of unlearning processes.
Appendix C Algorithm of the Frameworks
The procedure steps in the SplitWiper (cf. Sec.3) can be found from Algorithms 1 and 2, while that of its enhanced privacy-preserving option SplitWiper+ (cf. Sec.4) can be found from Algorithms 3 and 4.
Appendix D Experiments on Convergence
Establishing theoretical proof of convergence is challenging [13] for SplitWiper and SplitWiper+. Therefore, we rely on empirical experiments to assess convergence by tracking the number of nodes that change in each iteration. The experimental data indicates that the proportion of moved nodes steadily approaches zero within 30 epochs when using the CIFAR-10 dataset or a combination of CIFAR-10 and MNIST; for CIFAR-100, this occurs within 50 epochs.
Fig. 6 illustrates how test accuracy varies with increasing epochs for our frameworks compared to traditional SL frameworks, using the VGG and ResNet18 architectures across different datasets with heterogeneous label distributions among clients. The proposed SplitWiper and SplitWiper+ with various $\gamma$ and $\varepsilon$ consistently achieve higher accuracy than Vanilla SL and U-shaped SL across all datasets. It is also evident that test accuracy for all SL frameworks converges after a certain number of epochs, with CIFAR-100 requiring more epochs to reach convergence due to its larger label set. Since this trend is consistent across all scenarios and model architectures, we present one representative case with the highest degree of non-IID data distribution in Fig. 6 for analysis.
Table 5: Training and transmission time consumption (client-side) of different SL frameworks.
Data Distribution Dataset Scenario Client Type Training Time (s) Transmission Time (s)
U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+ U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+
$\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$ $\gamma=2$ $\gamma=3$ $\gamma=\Gamma_{p}$
Clients with same task Homo labels Hetero volumes CIFAR-10 Learning $C_{u}$$\star$ 74.36 66.03 2.23 9.73 15.85 9.08 216.40 261.63 9.04 18.08 27.13 18.08
$C_{o}$- 58.88 42.65 1.42 6.16 10.25 5.72 172.98 168.02 5.90 11.80 17.70 5.90
Unlearning $C_{u}$$\star$ 34.23 22.18 1.43 5.94 10.20 5.64 167.63 168.72 5.84 11.68 17.52 5.84
$C_{o}$- 57.48 42.89 0.00 0.00 0.00 0.00 172.95 168.42 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 79.88 74.68 1.44 6.31 10.70 5.88 324.58 328.02 8.24 16.48 24.73 24.72
$C_{o}$- 77.21 71.84 1.41 5.98 10.21 5.48 312.58 315.24 6.22 12.44 18.66 6.22
Unlearning $C_{u}$$\star$ 77.03 70.93 1.36 5.92 10.17 5.43 308.95 312.04 6.13 12.25 18.38 6.13
$C_{o}$- 77.15 71.56 0.00 0.00 0.00 0.00 309.45 312.04 0.00 0.00 0.00 0.00
Partially overlapping labels CIFAR-10 Learning $C_{u}$$\star$ 87.58 76.24 2.47 11.07 18.93 9.84 313.31 314.44 10.72 21.44 32.17 21.44
$C_{o}$- 42.99 38.59 1.28 5.53 9.43 5.47 162.61 157.22 5.34 10.68 16.02 10.42
Unlearning $C_{u}$$\star$ 67.88 53.51 1.74 7.72 13.31 6.41 219.54 220.83 7.52 15.04 22.56 15.04
$C_{o}$- 43.49 38.25 0.00 0.00 0.00 0.00 163.04 157.22 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 115.03 113.55 2.26 9.85 17.01 8.62 513.69 516.06 12.72 25.44 38.17 28.17
$C_{o}$- 68.54 60.44 1.08 4.43 7.56 4.19 285.63 287.24 5.28 10.56 15.84 14.32
Unlearning $C_{u}$$\star$ 112.48 110.02 2.16 9.53 16.44 8.04 498.07 500.07 12.40 24.80 37.21 14.40
$C_{o}$- 68.82 62.86 0.00 0.00 0.00 0.00 282.50 284.04 0.00 0.00 0.00 0.00
Hetero labels CIFAR-10 Learning $C_{u}$$\star$ 49.04 43.87 1.44 6.25 10.78 8.87 187.54 189.62 6.40 12.80 19.20 12.80
$C_{o}$- 50.31 44.34 1.46 6.28 10.78 7.37 187.53 189.63 6.40 12.80 19.21 12.82
Unlearning $C_{u}$$\star$ 27.80 22.18 0.73 3.12 5.39 5.57 93.77 96.01 3.20 6.41 9.57 9.60
$C_{o}$- 48.88 43.97 0.00 0.00 0.00 0.00 187.54 189.64 0.00 0.00 0.00 0.00
CIFAR-100 Learning $C_{u}$$\star$ 79.51 76.99 1.48 6.33 10.79 6.36 312.58 316.04 6.40 12.83 19.22 19.16
$C_{o}$- 79.74 77.02 1.50 6.35 10.83 5.78 312.60 316.03 6.42 12.80 19.22 11.21
Unlearning $C_{u}$$\star$ 78.88 76.66 1.44 6.00 10.26 5.79 296.96 300.04 6.08 12.16 18.24 12.16
$C_{o}$- 79.32 77.11 0.00 0.00 0.00 0.00 312.57 316.04 0.00 0.00 0.00 0.00
Clients with different tasks CIFAR-10 + MNIST Learning $C_{u}$$\star$ 93.61 88.92 2.95 13.50 31.64 13.08 381.11 384.05 12.80 25.61 37.29 24.76
$C_{o}$- 104.97 90.37 3.01 14.55 23.64 13.41 428.04 438.09 14.26 30.13 37.81 28.98
Unlearning $C_{u}$$\star$ 80.97 61.15 2.39 9.89 14.20 9.92 286.96 288.04 9.60 19.20 26.13 19.08
$C_{o}$- 103.79 90.35 0.00 0.00 0.00 0.00 428.09 437.94 0.00 0.00 0.00 0.00
Table 6: Test accuracy of different SL frameworks.
Data Distribution Dataset Scenario Label Type Accuracy (%)
U-shaped SL Vanilla SL SplitWiper ($\gamma=1$) SplitWiper+
$\gamma=2$ $\varepsilon=0.1$ $\gamma=2$ $\varepsilon=0.2$ $\gamma=3$ $\varepsilon=0.1$ $\gamma=3$ $\varepsilon=0.2$ $\gamma=\Gamma_{p}$ $\varepsilon=0.1$ $\gamma=\Gamma_{p}$ $\varepsilon=0.2$
Clients with same task Homo labels Hetero volumes CIFAR-10 Learning $Y_{u}$downarrow 95.40 92.70 99.80 98.10 98.70 97.40 97.90 99.40 99.70
$Y_{o}$uparrow 93.47 92.36 97.04 96.67 97.41 97.07 96.16 96.83 97.97
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 94.23 92.94 97.22 96.47 97.91 96.78 97.07 97.14 98.56
CIFAR-100 Learning $Y_{u}$downarrow 91.20 92.30 97.60 97.00 97.10 97.00 98.10 97.40 97.80
$Y_{o}$uparrow 82.01 81.77 85.39 83.75 84.17 82.53 83.44 85.27 86.08
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 82.79 82.02 86.86 83.51 84.95 82.71 84.23 86.19 87.17
Partially overlapping labels CIFAR-10 Learning $Y_{u}$downarrow 95.60 94.30 99.60 99.50 98.40 99.90 99.20 99.60 99.70
$Y_{o}$uparrow 94.03 92.91 97.30 96.34 97.26 95.80 96.94 96.79 97.78
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 95.08 92.35 97.72 97.61 97.19 96.11 96.64 97.81 98.07
CIFAR-100 Learning $Y_{u}$downarrow 91.90 91.50 98.00 97.70 98.00 97.60 97.60 98.00 98.10
$Y_{o}$uparrow 81.98 81.73 83.94 82.47 83.47 82.29 83.42 83.67 84.03
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 83.01 82.11 85.14 83.95 84.51 83.66 84.37 84.86 85.17
Hetero labels CIFAR-10 Learning $Y_{u}$downarrow 93.20 93.90 99.40 99.80 99.80 99.70 99.50 99.90 99.90
$Y_{o}$uparrow 93.04 93.67 98.52 97.15 97.92 95.74 96.92 98.26 98.10
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 93.93 94.13 98.60 96.98 98.53 97.06 97.61 98.59 98.14
CIFAR-100 Learning $Y_{u}$downarrow 90.30 92.10 99.90 99.70 98.90 98.70 99.00 99.00 99.10
$Y_{o}$uparrow 81.42 80.22 87.43 85.52 86.93 84.14 85.77 86.07 87.71
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 82.04 80.95 88.87 86.89 87.93 85.99 87.06 87.59 88.21
Clients with different tasks CIFAR-10 + MNIST Learning $Y_{u}$downarrow 95.80 95.60 99.90 97.20 98.60 96.80 96.80 97.30 97.40
$Y_{o}$uparrow 95.13 95.73 98.64 96.71 97.15 95.09 96.47 96.31 96.71
Unlearning $Y_{u}$downarrow 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
$Y_{o}$uparrow 95.12 95.74 99.07 96.89 97.01 95.33 96.34 96.56 96.72